The driver of a car travelling at 2m/s hits the brakes which provides a steady deceleration of 3m/s2 , how far does the vehicle travel before coming to rest?


1
Expert's answer
2022-03-23T07:15:29-0400

The distance travel by a car to the rest

"d=\frac{v_f^2-v_i^2}{2a}\\
=\frac{0^2-2^2}{2\times (-3)}=0.67\:\rm m"